Trace the curve y^3 = x^2 −1 , and state all the properties you use to trace it.
1
Expert's answer
2020-03-09T11:27:03-0400

The curve "y^3=x^2-1" coincides with the graph of the function "\quad \quad y(x)={ \left( { x }^{ 2 }-1 \right) }^{ \frac { 1 }{ 3 } }\quad \quad"

1) The function is defined on "\left( -\infty ,+\infty \right)" .

2) The function is continuous in the domain of definition, since it is elementary.

3) The function is even , so the graph of the function is symmetric about the axis 0Y.

4) "{ y }^{ ' }(x)=\frac { 2 }{ 3 } x{ \left( { x }^{ 2 }-1 \right) }^{ -\frac { 2 }{ 3 } },x\in \left( -\infty ,-1 \right) \cup \left( -1,1 \right) \cup \left( 1,+\infty \right)" .

5)By definition , the line "x=a" is the vertical tanget to the graph of the function at the point "(a,y(a))" if the function is continuous and "\lim _{ x\rightarrow a }{ \frac { y(x)-y(a) }{ x-a } } =\infty"

"y(1)=0, y(-1)=0" "\lim _{ x\rightarrow 1 }{ \frac { { \left( { x }^{ 2 }-1 \right) }^{ \frac { 1 }{ 3 } }-0 }{ x-1 } =\lim _{ x\rightarrow 1 }{ \frac { { \left( x+1 \right) }^{ \frac { 1 }{ 3 } } }{ { \left( x-1 \right) }^{ \frac { 2 }{ 3 } } } } =\infty } ," "\lim _{ x\rightarrow -1 }{ \frac { { \left( { x }^{ 2 }-1 \right) }^{ \frac { 1 }{ 3 } }-0 }{ x+1 } =\lim _{ x\rightarrow -1 }{ \frac { { \left( x-1 \right) }^{ \frac { 1 }{ 3 } } }{ { \left( x+1 \right) }^{ \frac { 2 }{ 3 } } } } =\infty } ." Thus, the function "y(x)" is not differentiable at the points "x=-1, x=1". The lines "x=-1,x=1" are the vertical tangents to the function graph.

6) The function increases in the intervals "\ \left( 0,1 \right) ,\ \left( 1,+\infty \right)" , since "y'(x)>0"

The function decreases in the intervals "\left( -\infty ,-1 \right) , \left( -1,0 \right)", since "y'(x)<0"

7) From 6) it follows that x=0 is a point of local and global minimum "\ { y }_{ min }=y(0)=-1" . "\lim _{ x\rightarrow +\infty }{ { \left( { x }^{ 2 }-1 \right) }^{ \frac { 1 }{ 3 } }=+\infty }" The range of function is "\left[- 1,+\infty \right)"

8) "\quad { y }^{ " }(x)=\\frac { 2 }{ 3 } { \\left[ { \\ \\left( { x }^{ 2 }-1 \\right) }^{ -\\frac { 2 }{ 3 } }-\\frac { 4 }{ 3 } { x }^{ 2 }{ \\left( { x }^{ 2 }-1 \\right) }^{ -\\frac { 5 }{ 3 } } \\right] }" "=-\frac { 2 }{ 9 } { \left( { x }^{ 2 }-1 \right) }^{ -\frac { 5 }{ 3 } }\left( { x }^{ 2 }+3 \right)"

"y"(x)>0" if "x\in [0,1)" , "y"(x)<0" if "x\in(1,+\infty)" . Hence function is convex down in "(0,1)" and convex up in "(1,+\infty)" .